OpenStudy (anonymous):

Prove the theorem that if limx→d h(x) = P and limx→d k(x) = Q and h(x) ≥ k(x) for all x in an open interval containing d, then P ≥ Q by using the formal definition of the limit. I know that it can be proved by contradiction assuming that P

OpenStudy (ikram002p):

ive seen this Qs before :/

OpenStudy (nincompoop):

because it is almost in every analysis/calculus textbook

OpenStudy (ikram002p):

aha but the prove want epsilon delta i think right ?

OpenStudy (anonymous):

Yeah I have to prove it with the formal definition of the limit epsilon deltas. not fun

OpenStudy (ikram002p):

aha i remember now .... i tried to prove to u once..

OpenStudy (nincompoop):

you know it says by definition, so I don't know why you're going to use by contradiction

OpenStudy (anonymous):

I can prove it any way but I thought it would be easier by contradiction. you said page 78 right? I cant see where you are looking

OpenStudy (nincompoop):

\[\lim_{x \rightarrow \Delta x}f(x)=\frac{ f(x+\Delta x)-f(x) }{ \Delta x }\] that is where the limit chapter starts

OpenStudy (anonymous):

I have to prove it with the epsilon delta definition of the limit so that one does not work.

OpenStudy (ikram002p):

wats ur cours name ??

OpenStudy (anonymous):

Math analysis the calc course after calc three

OpenStudy (ikram002p):

aha ok .. i thought that too its real analysis cours.. ok i found my old txt book at this sup(intoduction to real analysis )... i'll try to review n give u an answer.

OpenStudy (anonymous):

your great thanks!!

OpenStudy (anonymous):

Here is what I have right now: Prove the theorem that if limx→d h(x) = P and limx→d k(x) = Q and h(x) ≥ k(x) for all x in an open interval containing d, then P ≥ Q by using the formal definition of the limit Suppose that P<Q \[\left| h(x)-P \right|<e\] and\[\left| k(x)-Q \right|<e\] Let \[e=\frac{ (Q-P) }{ 2 }\] then there exist a delta (actually two of them lets choose the smaller) such that \[\left| x-d \right|<\delta\] now I know that there is a contradiction because as x gets closer to h(x) or k(x) it will get further away form the limits but I do not know how to show this???

OpenStudy (ikram002p):

ok i'll continue wat u got ... and see if we get contradicton ... now im not sure abt this i might be rong k ?? lim(h(x)-k(x)=P-Q |(h(x)-k(x))-(P_Q)|<e h(x)-k(x) converges to (P-Q) at d from hypo. & diff now assume f(x)=h(x)-g(x) ,f(x)>=0 given. |f(x)-(P-Q)|<e |-(P-Q)|<e......but |-(P-Q)|=|(P-Q)| |(P-Q)|<e -e<P-Q<e 0<P-Q+e<2e right ? now u got Q-P=2e apply it 0<P-Q+e<Q-P now u try to find a cont .... but i cant see it :/

OpenStudy (anonymous):

@ikram002p if I plug back for epsilon I get that p-q<q-p which is what I stated I need to use the fact that (x-d)<delta and show that if epsilon is equal to (q-p)/2 that i cannot find an x if P<Q. This one really sucks it is an statement that seems easy theoretically but I cannot prove it!! Thanks for all your help

OpenStudy (ikram002p):

ur wlc... even though i dint help u that much sry... good luck,,, and be carefull couse this course is triky trust me the limit is the esiest part :/ i hated that course .

OpenStudy (anonymous):

Thanks again!!

OpenStudy (anonymous):

I think I have it Prove the theorem that if limx→d h(x) = P and limx→d k(x) = Q and h(x) ≥ k(x) for all x in an open interval containing d, then P ≥ Q by using the formal definition of the limit Proof by contrdiction: Assume that Q>P \[\lim_{x \rightarrow d} h(x)-k(x) =P -Q\] therefore for any e>0, there exist delta>0 such that \[\left| (h(x)-k(x))-(P-Q) \right|<\epsilon \] whenever \[0<(x-d)<\delta\] Since Q-P> 0 by my assumption, lets take e=Q-P \[\left| (h(x)-k(x))-(P-Q) \right|Q-P\] Since a <= abs(a) for any number a \[(h(x)-f(x)) - (P-Q) < L=Q-P\] now reducing we get \[h(x)-k(x)<0\] or \[h)x)<k(x)\] This is a contridiction since it is stated that h(x) is greater then K(x) for all x in the open interval around d. Thus \[P \ge Q\] QED What do you think?

OpenStudy (anonymous):

@ikram002p

OpenStudy (ikram002p):

ok every thing sounds right untill this one " Since a <= abs(a) for any number a" now see i get a contradiction |k(x)-h(x)-(P-Q)|<Q-P ||k(x)-h(x)|-|(P-Q)||<=|k(x)-h(x)-(P-Q)|<Q-P ..but u know -|(P-Q)| sighn right ? so ||k(x)-h(x)|+(Q-P)|<=|k(x)-h(x)-(P-Q)|<Q-P and k(x)-h(x)>0 |k(x)-h(x)+(Q-P)|<Q-P=e which is a contradiction. any step u r confused in just ask :)